Solve the equation by completing the square.
X^2- 3x -8 =0

Answer by Ave(106)

x^2-3x = 8 add 8 to both sides so you can form the perfect square
(-3/2)^2 divide -3 by 2 and then square the answer for 3rd term
x^2-3x+9/4=8+9/4 add 9/4 to each side
(x-3/2)(x-3/2)= 41/4 reverse foil and combine like terms
(x-3/2)^2=41/4
x-3/2=ħsqrt(41/4) take the sqrt of both sides
x=3/2ħsqrt(41)/2 You can simplify more if you need to.